# Smooth Sailing: Lipschitz-Driven Uncertainty Quantification for Spatial Association

### v2

@@ Abstract
Estimating associations between spatial covariates and responses — rather than merely predicting responses — is central to environmental science, epidemiology, and economics. For instance, public health officials might be interested in whether air pollution has a strictly positive association with a health outcome, and the magnitude of any effect. Standard machine learning methods often provide accurate predictions but offer limited insight into covariate-response relationships. And we show that existing methods for constructing confidence (or credible) intervals for associations fail to provide nominal coverage in the face of model misspecification and distribution shift — despite both being essentially always present in spatial problems. We introduce a method that constructs valid frequentist confidence intervals for associations in spatial settings. Our method requires minimal assumptions beyond a form of spatial smoothness. In particular, we do not require model correctness or covariate overlap between training and target locations. Our approach is the first to guarantee nominal coverage in this setting and outperforms existing techniques in both real and simulated experiments.
@@ 1 Introduction
A priori, we might expect a nonlinear relationship between air pollution (as a covariate) and birth weight (as the response). A natural idea is to fit a flexible model — such as a (deep) Gaussian process [49, 13], transformer [64, 30], or XGBoost [10]. These methods can achieve high predictive accuracy, but the methods on their own often lack interpretability and do not immediately yield conclusions about covariate-response associations [52, 15]. Researchers have proposed a number of post hoc interpretability methods, such as Shapley values [56, 35], LIME [50], partial dependence plots [18] and Accumulated Local Effects plots [3]. These methods work directly with the fitted model and can be used to describe associations between individual covariates and model predictions. And because they are interpreting the fitted model (rather than the response itself), there is no need (or mechanism) to quantify uncertainties due to insufficient data. When the fitted model closely approximates the true response, one can view the output of these interpretability methods as describing the relationship between the covariates and response.
@@
It remains to choose an appropriate interpretable model (rather than post hoc method) and quantify its uncertainty. Buja et al., [7] demonstrated that (directly-fit, not post hoc) linear models can be used to interpretably summarize associations even in the face of potentially nonlinear relationships. The estimated parameters provide a best guess of association trends, and confidence intervals reflect uncertainty. And even in cases where data is sparse in the locations we are interested in, we might hope that we can learn simple linear associations between the covariate and response. These capabilities help explain why, despite known misspecification, linear models remain widely used. All of the applied studies cited above use linear regression. Moreover, Castro Torres and Akbaritabar, [9, Figure 1] surveyed quantitative methods in papers up to 2022 and found that over half of those in fields such as agricultural science, social sciences, and health sciences reported results from a linear model.
However, providing valid confidence intervals in this setting remains an open challenge. We are interested in the common scenario where the linear model is misspecified, and inference is needed at spatial locations that differ from those in the observed data. Under misspecification, Buja et al., [7] notes that the ordinary least squares (OLS) estimator depends on the covariate values, and classical confidence intervals are valid only at the observed (training) locations. The sandwich estimator [27, 71, 72, 7] offers valid intervals under misspecification, but only if all covariates are drawn from a single distribution. Generalized least squares (GLS) regression [2] is designed to handle spatial correlation in residuals [11, pp. 22–24], but it does not address the bias introduced by misspecification and distribution shift. Bayesian spatial models, such as those based on Gaussian processes, are also common, but their credible intervals tend to underestimate uncertainty in the presence of both misspecification and distribution shift [68, 39]. Finally, importance weighting methods [57] face two issues: (1) They assume one is able to debias the estimator, which need not be true, and (2) the validity of the confidence interval relies on perfectly estimated importance weights — an unrealistic assumption. We discuss these and related approaches in more detail in Appendix A.
@@
In what follows, we show through real and simulated experiments (Section 4) that all of the aforementioned methods can yield coverage far below the nominal level. Our principal contribution is to introduce the first method for constructing confidence intervals in spatial linear regression that guarantees frequentist coverage at the nominal level (a) under nonparametric assumptions on the data-generating distribution and (b) when the target locations differ from the observed ones. We are able to account for misspecification and covariate shift simultaneously by making more spatially appropriate assumptions than prior work. In particular, prior work relies on a (spatially inappropriate) assumption of independent and identically distribution data; we instead assume the response varies smoothly as a function of space. In our experiments, our method is the only one to consistently attain nominal coverage (or even come close).
@@ 2 Problem Setup
Data. Following the covariate-shift literature [4, 44, 12], we refer to our fully observed (training) data as the source data; likewise, we let target data denote the (test) locations and covariates where we do not observe the response but would like to understand the association between covariates and response. In particular, the source data consists of $N$ triplets $(S_{n},X_{n},Y_{n})_{n=1}^{N}$, with spatial location $S_{n}\in\mathcal{S}$, covariate $X_{n}\in\mathbb{R}^{P}$, and response $Y_{n}\in\mathbb{R}$. Here $\mathcal{S}$ represents geographic space; we assume $\mathcal{S}$ is a metric space with metric $d_{\mathcal{S}}$. We collect the source covariates in the matrix $X\in\mathbb{R}^{N\times P}$ and the source responses in the $N$-long column vector $Y$. The target data consists of $M$ pairs $(S^{\star}_{m},X^{\star}_{m})_{m=1}^{M}$, with $S^{\star}_{m}\in\mathcal{S}$, $X^{\star}_{m}\in\mathbb{R}^{P}$. The corresponding responses $Y^{\star}_{m}\in\mathbb{R},1\leq m\leq M$ are unobserved. We collect the target covariates in $X^{*}\in\mathbb{R}^{M\times P}$ and responses in column vector $Y^{*}\in\mathbb{R}^{M}$.
@@
where $\rho^{2}=\sigma^{2}e_{p}^{\mathrm{T}}(X^{\mathrm{T}}X)^{-1}e_{p}$, $z_{\alpha}$ is the $\alpha$ quantile of a standard normal distribution, and $e_{p}$ denotes the $P$-dimensional vector with a $1$ in entry $p$ and 00s elsewhere. Under correct specification, the confidence interval is valid: that is, it provides nominal coverage. E.g., a 95% confidence interval contains the true parameter at least 95% of the time under resampling. The noise variance $\sigma^{2}$ in $\rho$ used in LABEL:{eqn:ols-point-estimate-and-ci} is typically replaced with an estimate $\hat{\sigma}^{2}$.
Misspecified Spatial Setup: Data-Generating Process. In what follows, we assume the data is generated as $Y_{n}=f(X_{n},S_{n})+\epsilon_{n}$ and $Y^{\star}_{m}=f(X^{\star}_{m},S^{\star}_{m})+\epsilon^{\star}_{m},$ for a function $f$ that need not have a parametric form and with $\epsilon_{n},\epsilon^{\star}_{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for some (unknown) $\sigma^{2}>0$.
For ease of development, we will next make the simplifying assumption that source and target covariates are fixed functions of spatial location. Recall that all of the covariates in our examples (aerosol concentrations, proximity to highways, and air pollution) can be expected to vary spatially and be measured with minimal error. We similarly expect meteorological variables such as precipitation, humidity, and temperature to be reasonably captured by this assumption. [^1] Conversely demographic covariates may more reasonably be thought of as noisy functions of space, and further work is needed to handle the noisy case.
@@
Under Assumption 1, our data-generating process simplifies.
@@
As an aside, we note that one could instead treat the source and target locations as random draws from separate distributions, drop the conditioning on $S^{\star}_{m}$ (in Eq. 3), and rely on covariate-shift methods. Yet to the best of our knowledge these methods currently offer no finite-sample confidence intervals, leaving the question of valid confidence intervals unresolved. Moreover, in the spatial applications we study, locations are rarely i.i.d.; for example, United States Environmental Protection Agency (EPA) monitoring stations are placed strategically, and the targets of interest may be only a few — or sometimes just one — municipalities with missing data, so it is unclear whether there is a meaningful population representation. We therefore retain the conditional estimand in Eq. 3.
To solve the minimization in Eq. 3, it will be convenient to assume invertibility, as for OLS.
@@
With Assumptions 1 and 3, we can solve the minimization in Eq. 3 to find
@@
See Section E.1 for a derivation. To the best of our knowledge, the target-conditional estimand in Eq. 4 has not been previously proposed or analyzed in spatial linear regression.
In order to estimate $\mathbb{E}[Y^{\star}\mid S^{\star}]$ sufficiently well to in turn estimate $\theta^{\star}_{\text{OLS}}$ well, we need to make regularity assumptions. In classical and covariate-shift settings, these take the form of i.i.d. and bounded-density-ratio assumptions. Since we’ve seen that the classic assumptions are inappropriate in this spatial setting, we instead assume $f$ is shared across source and target (Assumption 2) and not varying so quickly in space as to be difficult to learn from limited data.
@@
As an illustration of how this assumption might be satisfied, consider $\chi$ an $L_{1}$-Lipschitz function of the spatial domain, and $f(s)=\beta^{\mathrm{T}}\chi(s)+g(s)$ with $g$ a fixed, $L_{2}$-Lipschitz function of $\mathcal{S}$. Then $f$ is a $(\|\beta\|_{2}L_{1}+L_{2})$-Lipschitz function of the spatial domain. A similar assumption was recently considered in Burt et al., [8] to derive a consistent estimator for prediction error in spatial settings.
@@ 3 Lipschitz-Driven Inference
We next provide a confidence interval for $\theta^{\star}_{\text{OLS},p}$, the $p$th coefficient of the target-conditional least squares estimand (Eq. 4). We support its validity with theory (in the present section) and experiments (in Section 4). To that end, we start by providing an efficiently-computable point estimate. We end by discussing the role and choice of the Lipschitz constant $L$.
Lipschitz-Driven Point Estimation. Since the target covariates are known, the key challenge in estimating Eq. 4 is estimating the unknown quantity $\mathbb{E}[Y^{\star}|S^{\star}]$.
@@
$\Psi_{mn}$ =$=\begin{cases}1&S_{n}=\textrm{closest source location to }S^{\star}_{m}\\ 0&\text{otherwise}\end{cases}.\textrm{ We break ties uniformly at random. }$. (5)
While this simple construction works well in our present experiments, we discuss the potential benefits of other constructions in Section B.1.
Lipschitz-Driven Confidence Intervals. We detail how to efficiently compute our proposed confidence interval for $\theta^{\star}_{\text{OLS},p}$ in Algorithm 1 (Appendix B). We prove its validity in Theorem 7 below. Before stating our result, we establish relevant notation and intuition for how our method works. First, we show the difference between our estimand and the estimator is normally distributed. Toward that goal, we start by writing $\theta^{\star}_{\text{OLS},p}-\hat{\theta}_{p}^{\Psi}=\sum\nolimits_{m=1}^{M}w_{m}f(S^{\star}_{m})-\sum\nolimits_{n=1}^{N}v^{\Psi}_{n}Y_{n},$ for $w:=e_{p}^{\mathrm{T}}\left(X^{\star\mathrm{T}}X^{\star}\right)^{-1}X^{\star\mathrm{T}}\in\mathbb{R}^{M}$ and $v^{\Psi}:=w\Psi\in\mathbb{R}^{N}$. By Assumption 2 and the previous equation,
@@
To bound the bias $b$, we use Assumption 4 to write
@@
where $\mathcal{F}_{L}$ is the space of $L$-Lipschitz functions from $\mathcal{S}\to\mathbb{R}$. In Section E.3, we show that it is possible to use Kantorovich-Rubinstein duality to restate the right side of Eq. 7 as a Wasserstein-1 distance between discrete measures. This alternative formulation is useful since it can be cast as a linear program [46, Chapter 3]; see Section E.3. Let $B$ denote the output of this linear program.
Given $B\geq|b|$, the following lemma (with proof in Section E.4) allows us to construct a confidence interval for $\theta^{\star}_{\text{OLS},p}$ centered on $\hat{\theta}_{p}^{\Psi}$. We discuss the benefits of this construction over alternative approaches in Section B.3.
@@
Let $b\in[-B,B]$, $\tilde{c}>0$, and $\alpha\in(0,1)$. Then the narrowest $1-\alpha$ confidence interval that is symmetric and valid for all $\mathcal{N}(b,\tilde{c}^{2})$ is of the form $[-B-\tilde{c}\Delta,B+\tilde{c}\Delta]$ where $\Delta$ is the solution of $\Phi\left(\Delta\right)-\Phi\left(-2B/\tilde{c}-\Delta\right)=1-\alpha$ with $\Phi$ the cumulative density function of a standard normal distribution. Also, the $\Delta$ satisfying this inequality is $\Delta\in[\Phi^{-1}(1-\alpha),\Phi^{-1}(1-\frac{\alpha}{2})]$.
The resulting confidence interval appears in Algorithm 1. We next establish its validity. So far, we have covered only the known $\sigma^{2}$ case. We handle the unknown $\sigma^{2}$ case after the following theorem.
@@
Suppose $(S^{\star}_{m},X^{\star}_{m},Y^{\star}_{m})_{m=1}^{M}$ and $(S_{n},X_{n},Y_{n})_{n=1}^{N}$ satisfy Assumptions 1, 4, 3 and 2 with known $\sigma^{2}$. Define the (random) interval $I^{\Psi}$ as in Algorithm 1 using the known value of $\sigma^{2}$. Then with probability at least $1-\alpha$, $\theta^{\star}_{\text{OLS},p}\in I^{\Psi}$. That is, $I^{\Psi}$ has coverage (conditional on the test locations) at least $1-\alpha$.
In Section E.2, we prove validity of our confidence interval for a generic choice of weight matrix $\Psi$. Theorem 7 is an immediate corollary of that result.
Consistent Estimation of the Noise Variance $\sigma^{2}$. Generally, the noise variance $\sigma^{2}$ is unknown, so we will substitute an estimate for $\sigma^{2}$ in the calculation of the confidence interval $I^{\Psi}$ (Step 5 in Algorithm 1). In Corollary 9 below, we show that the resulting confidence interval has asymptotically valid coverage. To that end, we first show that the estimator in Eq. 10 is consistent for $\sigma^{2}$.
@@
Suppose the spatial domain $\mathcal{S}=[-A,A]^{D}$ for some $A>0,D\in\mathbb{N}$. Take Assumptions 2 and 4. For any sequence of source spatial locations $(S_{n})_{n=1}^{\infty}$, take $\hat{\sigma}^{2}_{N}$ as in Eq. 10. Then $\hat{\sigma}^{2}_{N}\to\sigma^{2}$ in probability as $N\rightarrow\infty$.
See Section E.5 for a proof. For intuition, recall that the conditional expectation minimizes expected squared error over all functions. Since the conditional expectation is $L$-Lipschitz (Assumption 4),
@@
The empirical version of Eq. 8 is Eq. 10, which we show is a good estimate for $\sigma^{2}$ for large $N$.
When $\mathcal{S}$ is a subset of Euclidean space or a subset of the sphere, the minimization in Eq. 10 yields a quadratic program for computing $\hat{\sigma}^{2}_{N}$. We provide implementation details in Section B.5.
Given Proposition 8, it follows from Slutsky’s Lemma that the resulting confidence interval is asymptotically valid. We provide a formal statement below, and a proof in Section E.5

### v3

@@ Abstract
Estimating associations between spatial covariates and responses — rather than merely predicting responses — is central to environmental science, epidemiology, and economics. For instance, public health officials might be interested in whether air pollution has a strictly positive association with a health outcome, and the magnitude of any effect. Standard machine learning methods often provide accurate predictions but offer limited insight into covariate-response relationships. And we show that existing methods for constructing confidence (or credible) intervals for associations can fail to provide nominal coverage in the face of model misspecification and nonrandom locations — despite both being essentially always present in spatial problems. We introduce a method that constructs valid frequentist confidence intervals for associations in spatial settings. Our method requires minimal assumptions beyond a form of spatial smoothness and a homoskedastic Gaussian error assumption. In particular, we do not require model correctness or covariate overlap between training and target locations. Our approach is the first to guarantee nominal coverage in this setting and outperforms existing techniques in both real and simulated experiments. Our confidence intervals are valid in finite samples when the noise of the Gaussian error is known, and we provide an asymptotically consistent estimation procedure for this noise variance when it is unknown.
@@ 1 Introduction
Our goal in the present work is to provide valid and useful confidence intervals for an estimator of these associations when (i) the underlying model may be misspecified and (ii) inference is needed at spatial locations that may differ from those in the observed data. First we argue that existing methods do not already solve this problem. In particular, we make this argument in turn for modern flexible machine learning methods, linear methods, spatial regression methods, and debiasing approaches.
Flexible Machine Learning Methods. A priori, we might expect a nonlinear relationship between air pollution (as a covariate) and birth weight (as the response). A natural idea is to fit a flexible model — such as a (deep) Gaussian process [52, 15], transformer [70, 33], or XGBoost [10]. These methods can achieve high predictive accuracy, but the methods on their own often lack interpretability and do not immediately yield conclusions about covariate-response associations [57, 17]. Researchers have proposed a number of post hoc interpretability methods, such as Shapley values [61, 38], LIME [53], partial dependence plots [20] and Accumulated Local Effects plots [3]. These methods work directly with the fitted model and can be used to describe associations between individual covariates and model predictions. And because they are interpreting the fitted model (rather than the response itself), there is no need (or mechanism) to quantify uncertainties due to insufficient data. When the fitted model closely approximates the true response, one can view the output of these interpretability methods as describing the relationship between the covariates and response.
@@
Ordinary Least Squares. So it seems natural to choose an appropriate interpretable model (rather than post hoc method) and quantify its uncertainty. Buja et al., [7] demonstrated that (directly-fit, not post hoc) linear models can be used to interpretably summarize associations even in the face of potentially nonlinear relationships (i.e., under misspecification). This observation helps explain why all of the applied studies cited above use linear regression. [^1] In fact, Castro Torres and Akbaritabar, [9, Figure 1] surveyed quantitative methods in papers up to 2022 and found that over half of those in fields such as agricultural science, social sciences, and health sciences reported results from a linear model. Moreover, ordinary least squares (OLS) comes equipped with classical confidence intervals. However, when the linear model is misspecified, Buja et al., [7] notes that the OLS estimator depends on the covariate values, and it follows that classical confidence intervals are valid only at the observed (training) values. Since we’re interested in valid intervals at other locations, which will have different covariate values, we need another approach. The sandwich estimator [30, 77, 78, 7] offers valid intervals under misspecification, but only if all covariates are drawn from a single distribution, which isn’t the case in our setting where we want to draw inferences at different locations from where we observe data. Thus, even though OLS provides interpretable estimates, it does not solve our key challenge: constructing valid confidence intervals in the common scenario where both (i) the linear model is misspecified and (ii) inference is needed at spatial locations that may differ from those in the observed data.
Spatial Regression Methods. One might hope that a method specifically tailored for spatial regression could address these issues. Generalized least squares (GLS) regression [2] is designed to handle spatial correlation in residuals [13, pp. 22–24], but it does not address the bias introduced by misspecification and nonrandom locations. Bayesian spatial models, such as those based on Gaussian processes, are also common, but their credible intervals tend to underestimate uncertainty in the presence of both misspecification and nonrandom locations [74, 42].
Debiasing Approaches. Another natural idea is to construct a debiased estimator of the association, and then account for the variance of this debiasing procedure. Importance weighting methods from the covariate shift literature [62] pursue this goal by reweighting each observation according to an estimated density ratio between target and source covariate distributions, aiming to remove bias from distribution shift. Semiparametric inference in partially linear models (e.g., [55, 54, 11]) takes a related approach: it first fits a flexible model to capture variation explained by a “nuisance” parameter (in our case, the spatial location) and then constructs a debiased estimator of the target association by regressing out the variation explained by this nuisance, so that the remaining signal reflects the direct relationship between the covariate and the response. However, these methods face two key limitations. First, they assume that the debiasing step can perfectly remove bias, which need not be true (especially in finite samples). Second, their validity relies on the assumption that observations are drawn independently and identically. And this assumption rarely holds in spatial applications, where locations are fixed and often spread unevenly due to physical, logistical, or policy constraints. We discuss these and related approaches in more detail in Appendix˜A.
@@
Our Contribution. In what follows, we show through real and simulated experiments (Section˜4) that existing approaches can yield confidence intervals with coverage far below the nominal level. Our principal contribution is to introduce the first method for constructing confidence intervals in spatial associations that guarantees frequentist coverage at the nominal level even when the underlying model is misspecified and inference is required at fixed, nonrandom locations that differ from those observed. We are able to account for misspecification and nonrandom locations simultaneously by making more spatially appropriate assumptions than prior work. In particular, prior work relies on a (spatially inappropriate) assumption of independent and identically distribution data; we instead assume the response is a smooth function of space observed with homoskedastic Gaussian noise. When the variance of this noise is known, our confidence intervals are valid in finite samples. To address the common case where the variance is unknown, we provide an asymptotically consistent estimator for it. In our experiments, our method is the only one that consistently attains nominal coverage (or even comes close).
@@ 2 Problem Setup
Data. Following the covariate-shift literature [4, 47, 14], we refer to our fully observed (training) data as the source data; likewise, we let target data denote the (test) locations and covariates where we do not observe the response but would like to understand the association between covariates and response. In particular, the source data consists of $N$ triplets $(S_{n},X_{n},Y_{n})_{n=1}^{N}$, with spatial location $S_{n}\in\mathcal{S}$, covariate $X_{n}\in\mathbb{R}^{P}$, and response [^2] We propose possible extensions to multivariate responses in Appendix C. But we leave the multivariate-response case largely as an area for future work. $Y_{n}\in\mathbb{R}$. Here $\mathcal{S}$ represents geographic space; we assume $\mathcal{S}$ is a metric space with metric $d_{\mathcal{S}}$. We collect the source covariates in the matrix $X\in\mathbb{R}^{N\times P}$ and the source responses in the $N$-long column vector $Y$. The target data consists of $M$ pairs $(S^{\star}_{m},X^{\star}_{m})_{m=1}^{M}$, with $S^{\star}_{m}\in\mathcal{S}$, $X^{\star}_{m}\in\mathbb{R}^{P}$. The corresponding responses $Y^{\star}_{m}\in\mathbb{R},1\leq m\leq M$ are unobserved. We collect the target covariates in $X^{*}\in\mathbb{R}^{M\times P}$ and responses in column vector $Y^{*}\in\mathbb{R}^{M}$.
@@
where $\rho^{2}=\sigma^{2}e_{p}^{\mathrm{T}}(X^{\mathrm{T}}X)^{-1}e_{p}$, $z_{\alpha}$ is the $\alpha$-quantile of a standard normal distribution, and $e_{p}$ denotes the $P$-dimensional vector with a $1$ in entry $p$ and 0s elsewhere. Under correct specification, the confidence interval is valid: that is, it provides nominal coverage. E.g., a 95% confidence interval contains the true parameter at least 95% of the time under resampling. The noise variance $\sigma^{2}$ in $\rho$ used in LABEL:{eqn:ols-point-estimate-and-ci} is typically replaced with an estimate $\hat{\sigma}^{2}$.
Misspecified Spatial Setup: Data-Generating Process. In what follows, we assume the data is generated as $Y_{n}=g(X_{n},S_{n})+\epsilon_{n}$ and $Y^{\star}_{m}=g(X^{\star}_{m},S^{\star}_{m})+\epsilon^{\star}_{m},$ for a function $g$ that need not have a parametric form and with $\epsilon_{n},\epsilon^{\star}_{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for some (unknown) $\sigma^{2}>0$.
We assume that source and target covariates are fixed functions of spatial location. Recall that all of the covariates in our examples (aerosol concentrations, proximity to highways, and air pollution) can be expected to vary spatially and be measured with minimal error. We similarly expect meteorological variables such as precipitation, humidity, and temperature to be reasonably captured by this assumption. [^3] Conversely demographic covariates may more reasonably be thought of as noisy functions of space, and further work is needed to handle the noisy case.
@@
Under Assumption˜1, our data-generating process simplifies.
@@
We emphasize that we are still allowing the response to be a function of both the covariates and the spatial location. But we need not state the dependence on the covariates explicitly in Assumption˜2 due to Assumption˜1. Formally, let $g$ be the function satisfying $E[Y\mid X,S]=g(X,S)$. By Assumption˜1, $X=\chi(S)$, and so $E[Y\mid X,S]=g(\chi(S),S)$. Define $f:\mathcal{S}\rightarrow\mathbb{R}$ by $f(s)=g(\chi(S),S)$. In other words, the response can be a function of the covariates, but because the covariates are a fixed function of space, we can also write the response as a function of space alone.
@@
Before solving the minimization in Eq.˜3, we argue that covariate shift, as commonly defined, does not solve the problem of interest in this paper. Instead of using our estimand above, one might instead treat the source and target locations as random draws from separate distributions, drop the conditioning on $S^{\star}_{m}$ (in Eq.˜3), and rely on covariate-shift methods. Yet to the best of our knowledge these methods currently offer no finite-sample confidence intervals, leaving the question of valid confidence intervals unresolved. Moreover, in the spatial applications we study, locations are rarely i.i.d.: for example, the United States Environmental Protection Agency (EPA) places monitoring stations strategically, and the targets of interest may be only a few municipalities — or even a single one — with missing data. In such settings, it is unclear that a meaningful population-level distribution of locations exists. We therefore condition on the training and target locations in Eq.˜3, i.e. we treat the locations as deterministic. Alternatively, we can view our setting as a special case of covariate shift in which the training and target distributions are the respective fixed sets of locations, and therefore have disjoint support. However, under this view, standard covariate-shift estimators are inapplicable.
To solve the minimization in Eq.˜3, it will be convenient to assume invertibility, as for OLS.
@@
With Assumptions˜1 and 3, we can solve the minimization in Eq.˜3 to find
@@
See Section˜F.1 for a derivation. To the best of our knowledge, the target-conditional estimand in Eq.˜4 has not been previously proposed or analyzed in spatial linear regression.
In order to estimate $\mathbb{E}[Y^{\star}\mid S^{\star}]$ sufficiently well to in turn estimate $\theta^{\star}_{\text{OLS}}$ well, we need to make regularity assumptions. In classical and covariate-shift settings, these take the form of i.i.d. and bounded-density-ratio assumptions. Since we’ve seen that the classic assumptions are inappropriate in this spatial setting, we instead assume $f$ is shared across source and target (Assumption˜2) and not varying so quickly in space as to be difficult to learn from limited data.
@@
As an illustration of how this assumption might be satisfied, consider $\chi$ an $L_{1}$-Lipschitz function of the spatial domain, and $f(s)=\beta^{\mathrm{T}}\chi(s)+h(s)$ with $h$ a fixed, $L_{2}$-Lipschitz function of $\mathcal{S}$. Then $f$ is a $(\|\beta\|_{2}L_{1}+L_{2})$-Lipschitz function of the spatial domain. A similar assumption was recently considered in Burt et al., [8] to derive a consistent estimator for prediction error in spatial settings.
Recall (from the discussion of $g$ vs. $f$ after Assumption˜2) that we ultimately allow our response to be a function of both the covariates and the spatial location. But, using our assumption that $f$ covers both dependencies (Assumption˜2), we impose our smoothness restrictions on the average response as a function of space alone. We think it is often easier to reason about smoothness in physical space, rather than in a (potentially high-dimensional) space of covariates.
@@ 3 Lipschitz-Driven Inference
We next provide a confidence interval for $\theta^{\star}_{\text{OLS},p}$, the $p$th coefficient of the target-conditional least squares estimand (Eq.˜4). We support its validity with theory (in the present section) and experiments (in Section˜4). To that end, we start by providing an efficiently-computable point estimate. We end by discussing the role and choice of the Lipschitz constant $L$.
Lipschitz-Driven Point Estimation. Since the target covariates are known, the key challenge in estimating Eq.˜4 is estimating the unknown quantity $\mathbb{E}[Y^{\star}|S^{\star}]$.
@@
[…]
While this simple construction works well in our present experiments, we discuss the potential benefits of other constructions in Section˜B.1. This point estimation approach is closely related to KNN imputation [66], used for missing data. But KNN imputation does not account for repeated use of training responses or bias in estimation due to imputation, problems we address in the next section. We provide a more complete discussion of KNN imputation in Appendix˜A.
Lipschitz-Driven Confidence Intervals. We detail how to efficiently compute our proposed confidence interval for $\theta^{\star}_{\text{OLS},p}$ in Algorithm˜1 (Appendix˜B). We prove its validity in Theorem˜7 below. Before stating our result, we establish relevant notation and intuition for how our method works. First, we show the difference between our estimand and the estimator is normally distributed. Toward that goal, we start by writing $\theta^{\star}_{\text{OLS},p}-\hat{\theta}_{p}^{\Psi}=\sum\nolimits_{m=1}^{M}w_{m}f(S^{\star}_{m})-\sum\nolimits_{n=1}^{N}v^{\Psi}_{n}Y_{n},$ for $w:=e_{p}^{\mathrm{T}}\left(X^{\star\mathrm{T}}X^{\star}\right)^{-1}X^{\star\mathrm{T}}\in\mathbb{R}^{M}$ and $v^{\Psi}:=w\Psi\in\mathbb{R}^{N}$. By Assumption˜2 and the previous equation,
@@
To bound the bias $b$, we use Assumption˜4 to write
@@
where $\mathcal{F}_{L}$ is the space of $L$-Lipschitz functions from $\mathcal{S}\to\mathbb{R}$. In Section˜F.3, we show that it is possible to use Kantorovich–Rubinstein duality to restate the right side of Eq.˜7 as a Wasserstein-1 distance between discrete measures. This alternative formulation is useful since it can be cast as a linear program [49, Chapter 3]; see Section˜F.3. Let $B$ denote the output of this linear program.
Given $B\geq|b|$, the following lemma (with proof in Section˜F.4) allows us to construct a confidence interval for $\theta^{\star}_{\text{OLS},p}$ centered on $\hat{\theta}_{p}^{\Psi}$. We discuss the benefits of this construction over alternative approaches in Section˜B.5.
@@
Let $b\in[-B,B]$, $\tilde{c}>0$, and $\alpha\in(0,1)$. Then the narrowest $1-\alpha$ confidence interval that is symmetric and valid for all $\mathcal{N}(b,\tilde{c}^{2})$ is of the form $B-\tilde{c}\Delta,B+\tilde{c}\Delta]$ where $\Delta$ is the solution of $\Phi\left(\Delta\right)-\Phi\left(-2B/\tilde{c}-\Delta\right)=1-\alpha$ with $\Phi$ the cumulative density function of a standard normal distribution. Also, the $\Delta$ satisfying this inequality is $\Delta\in[\Phi^{-1}(1-\alpha),\Phi^{-1}(1-\frac{\alpha}{2})]$.
The resulting confidence interval appears in Algorithm˜1. We next establish its validity. So far, we have covered only the known $\sigma^{2}$ case. We handle the unknown $\sigma^{2}$ case after the following theorem.
@@
Suppose $(S^{\star}_{m},X^{\star}_{m},Y^{\star}_{m})_{m=1}^{M}$ and $(S_{n},X_{n},Y_{n})_{n=1}^{N}$ satisfy Assumptions˜1, 4, 3 and 2 with known $\sigma^{2}$. Define the (random) interval $I^{\Psi}$ as in Algorithm˜1 using the known value of $\sigma^{2}$. Then with probability at least $1-\alpha$, $\theta^{\star}_{\text{OLS},p}\in I^{\Psi}$. That is, $I^{\Psi}$ has coverage (conditional on the test locations) at least $1-\alpha$.
In Section˜F.2, we prove validity of our confidence interval for a generic choice of weight matrix $\Psi$. Theorem˜7 is an immediate corollary of that result.
Consistent Estimation of the Noise Variance $\sigma^{2}$. Generally, the noise variance $\sigma^{2}$ is unknown, so we will substitute an estimate for $\sigma^{2}$ in the calculation of the confidence interval $I^{\Psi}$ (Step 5 in Algorithm˜1). In Corollary˜9 below, we show that the resulting confidence interval has asymptotically valid coverage. To that end, we first show that the estimator in Eq.˜10 is consistent for $\sigma^{2}$.
@@
Suppose the spatial domain $\mathcal{S}=[-A,A]^{D}$ for some $A>0,D\in\mathbb{N}$. Take Assumptions˜2 and 4. For any sequence of source spatial locations $(S_{n})_{n=1}^{\infty}$, take $\hat{\sigma}^{2}_{N}$ as in Eq.˜10. Then $\hat{\sigma}^{2}_{N}\to\sigma^{2}$ in probability as $N\rightarrow\infty$.
See Section˜F.5 for a proof. For intuition, recall that the conditional expectation minimizes expected squared error over all functions. Since the conditional expectation is $L$-Lipschitz (Assumption˜4),
@@
The empirical version of Eq.˜8 is Eq.˜10, which we show is a good estimate for $\sigma^{2}$ for large $N$.
When $\mathcal{S}$ is a subset of Euclidean space or a subset of the sphere, the minimization in Eq.˜10 yields a quadratic program for computing $\hat{\sigma}^{2}_{N}$. We provide implementation details in Section˜B.7.
Given Proposition˜8, it follows from Slutsky’s Lemma that the resulting confidence interval is asymptotically valid. We provide a formal statement below, and a proof in Section˜F.5